# SimBa: Simplicity Bias for Scaling Up Parameters in Deep Reinforcement Learning

### v1

@@ 3.2 Deep Reinforcement Learning
Despite these advances, scaling up network architectures in deep RL remains underexplored. While larger models improved performance in supervised learning by leveraging simplicity bias, this principle has not been fully explored in RL. Several recent studies have attempted to scale up network sizes—through ensembling (Chen et al., 2021; Obando-Ceron et al., 2024), widening (Schwarzer et al., 2023), and deepening networks (Bjorck et al., 2021; Nauman et al., 2024). However, they often rely on computationally intensive layers like spectral normalization (Bjorck et al., 2021) or require sophisticated training protocols (Nauman et al., 2024), limiting their applicability.
@@ 4 SimBa
This section introduces the SimBa network, an architecture designed to embed simplicity bias into deep RL. The architecture comprises Running Statistics Normalization, Residual Feedforward Blocks, and Post-Layer Normalization. By amplifying the simplicity bias, SimBa allows the model to avoid overfitting for highly overparameterized configurations.
@@
At each block, $l\in\{1,...,L\}$, the input $\mathbf{x}_{t}^{l}$ passes through a pre-layer normalization residual feedforward block, introducing simplicity bias by allowing a direct linear pathway from input to output. This direct linear pathway enables the network to pass the input unchanged unless non-linear transformations are necessary. Each block is defined as:
@@ 5.1 Architectural Components
Figure 4: Component Analysis. (a) Simplicity bias scores estimated via Fourier analysis. Mean and 95% CI are computed over 100 random initializations. (b) Average return in DMC-Hard at 1M steps. Mean and 95% CI over 10 seeds, using the SAC algorithm. Stronger simplicity bias correlates with higher returns for overparameterized networks.
@@ 5.2 Comparison with Other Architectures
To assess SimBa’s scalability, we compared it to BroNet (Nauman et al., 2024), SpectralNet (Bjorck et al., 2021), and MLP. Detailed descriptions of each architecture are provided in Appendix B.
The key differences between SimBa and other architectures lie in the placement of components that promote simplicity bias. SimBa maintains a direct linear residual pathway from input to output, applying non-linearity exclusively through residual connections. In contrast, other architectures introduce non-linearities within the input-to-output pathway, increasing functional complexity. Additionally, SimBa employs post-layer normalization to ensure consistent activation scales across all hidden dimensions, reducing variance in policy and value predictions. Conversely, other architectures omit normalization before the output layer, which can lead to high variance in their predictions.
To ensure that performance differences were attributable to architectural design choices rather than varying input normalization strategies, we uniformly applied RSNorm as the input normalization method across all models. Additionally, our investigation primarily focused on scaling the critic network’s hidden dimension, as scaling the actor network showed limited benefits (see Section7.2).
As illustrated in Figure 5.(a), SimBa achieves the highest simplicity bias score compared to the experimented architectures, demonstrating its strong preference for simpler solutions. In Figure 5.(b), while the MLP failed to scale with increasing network size, BroNet, SpectralNet, and SimBa all showed performance improvements as their networks scaled up.
Importantly, the scalability of each architecture was strongly correlated with its simplicity bias score, where a higher simplicity bias led to better scalability. SimBa demonstrated the best scalability, supporting the hypothesis that simplicity bias is a key factor in scaling deep RL models.
Figure 5: Architecture Comparison. (a) SimBa consistently exhibits a higher simplicity bias score. (b) SimBa demonstrates superior scaling performance in terms of average return for DMC-Hard compared to the other architectures. The results are from 5 random seeds.
@@ 6 Experiments
This section evaluates SimBa’s applicability across various deep RL algorithms and environments. For each baseline, we either use the authors’ reported results or run experiments using their recommended hyperparameters. Detailed descriptions of the environments used in our evaluations are provided in Appendix F. In addition, we also include learning curves and final performance for each task in Appendix H and I, along with hyperparameters used in our experiments in Appendix G.
@@ 7.1 Observation Normalization
A key factor in SimBa’s success is using RSNorm for observation normalization. To validate its effectiveness, we compare 5 alternative methods: (i) Layer Normalization (Ba et al., 2016); (ii) Batch Normalization (Ioffe & Szegedy, 2015); (iii) Env Wrapper RSNorm, which tracks running statistics for each dimension, normalizes observations upon receiving from the environment, and stores them in the replay buffer; (iv) Initial N Steps: fixed statistics derived from initially collected transition samples, where we used $N=5,000$; and (v) Oracle Statistics, which rely on pre-computed statistics from previously collected expert data.
@@ 8 Lessons and Opportunities
Lessons. Historically, deep RL has suffered from overfitting, necessitating sophisticated training protocols and implementation tricks to mitigate these issues (Hessel et al., 2018; Fujimoto et al., 2023). These complexities serve as significant barriers for practitioners with limited resources, hindering the effective usage of deep RL methods. In this paper, we improved performance by solely modifying the network architecture while keeping the underlying algorithms unchanged. This approach simplifies the implementation of SimBa, making it easy to adopt. By incorporating simplicity bias through architectural design and scaling up the number of parameters, the network converges to simpler, generalizable functions, matching or surpassing state-of-the-art deep RL methods. We also believe that SimBa follows the insights from Richard Sutton’s Bitter Lesson (Sutton, 2019): although task-specific designs can yield immediate improvements, an approach that scales effectively with increased computation may provide more sustainable benefits over the long term.
Opportunities. Our exploration of simplicity bias has primarily concentrated on the network architecture; however, optimization techniques are equally vital. Strategies such as dropout (Hiraoka et al., 2021), data augmentation (Kostrikov et al., 2020), and diverse optimization algorithms (Foret et al., 2020) can further enhance convergence to simpler functions during training. Integrating these techniques with continued parameter scaling presents a promising avenue for future research in deep RL. Furthermore, while our focus has been on the basic model-free algorithm SAC, there exists considerable empirical success with model-based algorithms (Schwarzer et al., 2020; Hansen et al., 2023; Hafner et al., 2023). Preliminary investigations into applying SimBa to model-based RL, specifically TD-MPC2 (Hansen et al., 2023), have shown promise. We encourage the research community to collaboratively pursue these avenues to advance deep RL architectures and facilitate their successful application in real-world scenarios.
@@ Acknowledgement
We would like to express our gratitude to Craig Sherstan, Bram Grooten, Hanseul Cho, Kyungmin Lee, and Hawon Jeong for their helpful discussions and insightful feedback on this paper.
@@ Appendix B Measuring Simplicity Bias
Appendix C Plasticity Analysis
@@
To assess how different architectures impact plasticity, we conducted experiments in DMC-Hard. We compared Soft Actor-Critic (SAC) implemented with a standard MLP architecture against SAC with the SimBa architecture. Each configuration was evaluated across 5 random seeds.
Figure 15: Plasticity Metrics Comparison. Average episode return for the DMC-Hard environment. Metrics compared include dormant ratio, s-rank, and feature norm. Higher dormant ratio and feature norm, along with lower s-rank, indicate a greater loss of plasticity.
Figure 15 shows that SAC with the MLP architecture exhibits a high dormant ratio, low s-rank, and large feature norms. These metrics indicate a significant loss of plasticity. In contrast, the SimBa architecture maintains lower dormant ratios, higher s-rank values, and balanced feature norms. This ensures more active neurons, diverse and rich feature representations, and stable activations.
These results demonstrate that SimBa effectively preserves plasticity, avoiding the degenerative trends seen with the MLP-based SAC. It will be interesting for future work to understand how simplicity bias and architectural choices influence network plasticity.
Appendix D Comparison to Existing Architectures
Here we provide a more in-depth discussion related to the architectural difference between SimBa, BroNet, and SpectralNet. As shown in Figure 16, SimBa differs from BroNet in three key aspects: (i) the inclusion of RSNorm, (ii) the implementation of pre layer-normalization, (iii) the utilization of a linear residual pathway, (iv) and the inclusion of a post-layer normalization layer. Similarly, compared to SpectralNet, SimBa incorporates RSNorm, employs a linear residual pathway, and leverages spectral normalization differently.
@@ Appendix D Comparison to Existing Architectures
Appendix E Computational Resources
@@
Appendix F Environment Details
@@
F.1 DeepMind Control Suite
@@ F.4 Craftax
Appendix G Hyperparameters
G.1 Off-Policy
@@ G.3 Unsupervised RL
Appendix H Learning Curve
@@
Appendix I Full Result

### v2

@@ 3.2 Deep Reinforcement Learning
Despite these advances, scaling up network architectures in deep RL remains underexplored, especially by leveraging the simplicity bias principle. Several recent studies have attempted to scale up network sizes—through ensembling (Chen et al., 2021; Obando-Ceron et al., 2024), widening (Schwarzer et al., 2023), residual connections (Espeholt et al., 2018) and deepening networks (Bjorck et al., 2021; Nauman et al., 2024). However, they often rely on computationally intensive layers such as spectral normalization (Bjorck et al., 2021) or require sophisticated training protocols (Nauman et al., 2024), limiting their applicability.
@@ 4 SimBa
This section introduces the SimBa network, an architecture designed to embed simplicity bias into deep RL. The architecture is composed of Running Statistics Normalization, Residual Feedforward Blocks, and Post-Layer Normalization. By amplifying the simplicity bias, SimBa allows the model to avoid overfitting for highly overparameterized configurations.
@@
At each block, $l\in\{1,...,L\}$, the input $\mathbf{x}_{t}^{l}$ passes through a pre-layer normalization residual feedforward block, introducing simplicity bias by allowing a direct linear pathway from the input to the output. This direct linear pathway enables the network to pass on the input unchanged throughout the entire network unless non-linear transformations are necessary. Each block is defined as:
@@ 5.1 Architectural Components
Figure 4: Component Analysis. (a) Simplicity bias scores estimated via Fourier analysis. Mean and 95% CI are computed over 100 random initializations. (b) Average return in DMC-Hard for 1M steps. Mean and 95% CI over 10 seeds, using SAC. Stronger simplicity bias correlates with higher returns for overparameterized networks.
@@ 5.2 Comparison with Other Architectures
To assess SimBa’s scalability, we compared it to BroNet (Nauman et al., 2024), SpectralNet (Bjorck et al., 2021), and MLP. Detailed descriptions of each architecture are provided in Appendix F.
The key differences between SimBa and other architectures lie in the placement of components that promote simplicity bias. SimBa maintains a direct linear residual pathway from the input to the output, applying non-linearity exclusively through residual connections. In contrast, other architectures introduce non-linearities within the input-to-output pathway, increasing functional complexity. Additionally, SimBa employs post-layer normalization to ensure consistent activation scales across all hidden dimensions, reducing variance in policy and value predictions. Conversely, other architectures omit normalization before the output layer, which can lead to high variance in their predictions.
To ensure that performance differences were attributable to architectural design choices rather than varying input normalization, we uniformly applied RSNorm as the input normalization across all models. Our investigation focused on scaling up the critic network’s parameters by increasing the hidden dimension, as scaling up the actor network showed limited benefits (see Section 7.2).
As illustrated in Figure 5.(a), SimBa achieves the highest simplicity bias score compared to the experimented architectures, demonstrating its strong preference for simpler solutions. In Figure 5.(b), while the MLP failed to scale with increasing network size, BroNet, SpectralNet, and SimBa all showed performance improvements.
Importantly, the scalability of each architecture is correlated with its simplicity bias score, with a higher simplicity bias leading to better scalability. SimBa demonstrated the best scalability, supporting the hypothesis that simplicity bias plays a key role in scaling deep RL. Further analysis on this correlation is provided in Appendix D.
Figure 5: Architecture Comparison. (a) SimBa consistently exhibits a higher simplicity bias score. (b) SimBa demonstrates strong scaling performance in terms of average return for DMC-Hard compared to the other architectures. The results are from 5 random seeds.
@@ 6 Experiments
This section evaluates SimBa’s applicability across various deep RL algorithms and environments. For each baseline, we either use the authors’ reported results or run experiments using their recommended hyperparameters. Detailed descriptions of the environments used in our evaluations are provided in Appendix H. In addition, we also include learning curves and final performance for each task in Appendix J and K, along with hyperparameters used in our experiments in Appendix I.
@@ 7.1 Observation Normalization
A key factor in SimBa’s success is using RSNorm for observation normalization. To validate its effectiveness, we compare 6 alternative methods: (i) LayerNorm (Ba et al., 2016); (ii) RMSNorm (Zhang & Sennrich, 2019); (iii) BatchNorm (Ioffe & Szegedy, 2015); (iv) Env Wrapper RSNorm, which tracks running statistics for each dimension, normalizes observations upon receiving from the environment, and stores them in the replay buffer; (v) Initial N Steps, which fixes the statistics derived from the initially collected transition samples, where we used $N=5,000$; and (vi) Oracle Statistics, which rely on pre-computed statistics from previously collected expert data.
@@ 8 Lessons and Opportunities
Lessons. Deep reinforcement learning has historically struggled with overfitting, requiring complex training protocols and tricks to mitigate these issues (Hessel et al., 2018; Fujimoto et al., 2023). These complexities can hinder practitioners with limited resources. In this paper, we improved performance solely by modifying the network architecture while keeping the underlying algorithms unchanged, simplifying SimBa’s implementation and making it easy to adopt. By incorporating simplicity bias through architectural design and scaling up parameters, our network converges to simpler, generalizable functions, matching or surpassing state-of-the-art methods. This aligns with Richard Sutton’s Bitter Lesson (Sutton, 2019): while task-specific designs may offer immediate gains, scalable approaches provide more sustainable long-term benefits.
Opportunities. While our exploration has focused on network architecture, optimization techniques such as dropout (Hiraoka et al., 2021), data augmentation (Kostrikov et al., 2020), and advanced optimization algorithms (Foret et al., 2020) are also crucial for promoting convergence to simpler functions. Extending our insights to vision-based RL presents a promising direction; specifically, designing convolutional architectures guided by simplicity bias principles could substantially enhance learning efficiency in visual tasks. Furthermore, integrating SimBa with multi-task RL may improve performance across diverse tasks by facilitating the learning of shared, simple, and generalizable representations across tasks. Although our current work centers on model-free RL algorithms, model-based approaches have also demonstrated considerable success (Schwarzer et al., 2020; Hansen et al., 2023; Hafner et al., 2023). Our preliminary investigations applying SimBa to model-based RL, specifically TD-MPC2 (Hansen et al., 2023), have yielded promising results that warrant further exploration. We encourage the research community to pursue these directions to advance deep RL architectures and facilitate their successful application in real-world scenarios.
@@ Acknowledgement
This work was supported by the Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)) and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2025-00555621). We would also like to express our gratitude to Craig Sherstan, Bram Grooten, Hanseul Cho, Kyungmin Lee, and Hawon Jeong for their invaluable discussions and insightful feedback, which have significantly contributed to this work.
@@ Appendix B Measuring Simplicity Bias
Appendix C Additional Ablations
C.1 Architectural Ablation
Experimental Setup. Following the architectural component analysis in Section 5.1, we further conducted ablation studies by removing each key component from SimBa to assess its impact on the simplicity score and average return. We performed these experiments on DMC-Hard using the same training setup as in Section 5.1, training for 1M environment steps.
Results. As shown in Figure 15.(a), we observe that, except for RSNorm, removing key components of SimBa—such as residual connections, pre-layer normalization, or post-layer normalization—leads to a reduction in the simplicity bias score. This indicates that the functional representation at initialization becomes biased toward more complex functions when these components are absent. This finding further emphasizes the critical role of each component in promoting simplicity bias within SimBa’s design. Additionally, Figure 15.(b) demonstrates that removing any key component significantly impacts SimBa’s performance, underscoring the necessity of each architectural component.
Figure 15: Additional Component Analysis. (a) Simplicity bias scores estimated via Fourier analysis. Mean and 95% CI are computed over 100 random initializations. (b) Average return in DMC-Hard for 1M steps. Mean and 95% CI over 10 seeds, using SAC.
Regarding RSNorm, although its absence significantly impacts the agent’s performance negatively, the overall simplicity bias score of the architecture increases rather than decreases. We speculate that this discrepancy arises because our Fourier Transform based complexity measure relies on evaluating the network’s outputs over a uniformly distributed input space. RSNorm is designed to normalize activations based on the statistics of non-uniform or dynamically changing input distributions, excelling with inputs that exhibit variance and complexity not captured by a uniform distribution. Therefore, since our simplicity analysis measure is based on a uniform input distribution, it may not fully capture RSNorm’s practical contribution to promoting simplicity bias.
C.2 Extended Training
Experimental Setup. We assess SimBa’s robustness by evaluating its performance on the DeepMind Control Suite (DMC) Hard tasks over an extended training period of 5 million timesteps. We compared three methods: SAC (Haarnoja et al., 2018), TD-MPC2 (Hansen et al., 2023), and SAC with SimBa. All models are trained under identical hyperparameters and configurations as in our main experiments to ensure a fair comparison.
Figure 16: Per task learning curve for DMC Hard. Mean and 95% CI are computed over 5 seeds for SimBa, 3 seeds for TD-MPC2 and SAC.
Results. Figure 16 presents the learning curve of each method for 5 million timesteps. The learning curves indicate that SimBa + SAC not only learns faster but also maintains more stable and consistent performance improvements over time, demonstrating better sample efficiency and reduced variance compared to the baselines.
Appendix D Correlation between Simplicity Bias and Performance
To investigate the relationship between simplicity bias score and performance in deep reinforcement learning, we computed the correlation between simplicity bias scores and average returns for DMC-Hard across various architectures evaluated in our study. Our analysis includes 12 architectures: four baselines architectures-MLP, SimBa, BroNet (Nauman et al., 2024), and SpectralNet (Bjorck et al., 2021)-along with eight SimBa variants. These variants consists of four variants with component additions (Section 5.1) and four variants with component reductions (Appendix C.1), representing a gradual transition from MLP to SimBa.
As discussed in Appendix F, the current simplicity bias measure may not accurately capture the influence of RSNorm. To address this limitation, we separately evaluated the correlation for architectures incorporating RSNorm. All twelve architectures have nearly identical parameter counts, approximately 4.5 million parameters each, with a maximum variation of 1%. Performance was assessed on the DMC-Hard benchmark, trained for one million steps.
Results. As illustrated in Figure 17(a), architectures employing RSNorm exhibit a Pearson correlation coefficient of $0.79$, indicating a strong positive correlation between simplicity bias score and average return in scaled architectures. In contrast, Figure 17(b) shows a Pearson correlation coefficient of $0.54$, representing a moderate correlation across all architectures. We attribute the lower correlation in the broader set to the current limitations of the simplicity bias measure. We believe these findings support our hypothesis that higher simplicity bias scores are associated with better performance in deep reinforcement learning with scaled-up architectures.
Figure 17: Correlation Analysis. Pearson correlation between simplicity bias score and average return on DMC-Hard. (a) Architectures incorporating RSNorm show a strong correlation, $\rho=0.79$. (b) All evaluated architectures exhibit a moderate correlation, $\rho=0.54$.
Appendix E Plasticity Analysis
@@ E.1 Metrics
E.2 Results
To evaluate the impact of different architectures on network plasticity, we compared the Soft Actor-Critic (SAC) algorithm implemented with a standard Multi-Layer Perceptron (MLP) against SAC with the SimBa architecture on the DMC-Easy & Medium and DMC-Hard tasks. Additionally, we conducted an ablation study to analyze how each component of SimBa contributes to plasticity measures. Each configuration was evaluated across five random seeds.
Figure 18: Plasticity Analysis of SimBa versus MLP. Metrics include dormant ratio, s-rank, and feature norm. Figures (a) and (b) represent DMC-Easy & Medium and DMC-Hard environments, respectively. A higher dormant ratio and feature norm, along with a lower s-rank, indicate a greater loss of plasticity.
As illustrated in Figures 18.(a) and (b), SAC with the MLP architecture exhibits high dormant ratios and large feature norms across both DMC-Easy & Medium and DMC-Hard tasks, indicating significant loss of plasticity. While MLP achieves higher s-rank values than SimBa in the DMC-Easy & Medium tasks, SimBa outperforms MLP in s-rank for the DMC-Hard tasks. Importantly, SimBa consistently maintains lower dormant ratios, balanced feature norms, and overall higher s-rank values in more complex environments.
These findings indicate that SimBa effectively preserves plasticity, avoiding the degenerative trends observed with the MLP-based SAC.
Figure 19: Plasticity Analysis of SimBa Components. Metrics include dormant ratio, s-rank, and feature norm. Figures (a) and (b) represent DMC-Easy & Medium and DMC-Hard environments, respectively. A higher dormant ratio and feature norm, along with a lower s-rank, indicate a greater loss of plasticity.
Furthermore, for the component ablation study (Figure 19), removing each component of SimBa individually shows diminished average returns. However, ablating each component does not significantly reduce plasticity measures, with only post-layer normalization showing an improved feature norm. This suggests that while individual components may not significantly impact plasticity measures with their sole use, their integrated use within the SimBa architecture is crucial for effectively mitigating plasticity loss.
Appendix F Comparison to Existing Architectures
Here we provide a more in-depth discussion related to the architectural difference between SimBa, BroNet, and SpectralNet. As shown in Figure 20, SimBa differs from BroNet in three key aspects: (i) the inclusion of RSNorm, (ii) the implementation of pre-layer normalization, (iii) the utilization of a linear residual pathway, (iv) and the inclusion of a post-layer normalization layer. Similarly, compared to SpectralNet, SimBa incorporates RSNorm, employs a linear residual pathway, and leverages spectral normalization differently.
@@ Appendix F Comparison to Existing Architectures
Appendix G Computational Resources
@@
Appendix H Environment Details
@@
H.1 DeepMind Control Suite
@@ H.4 Craftax
Appendix I Hyperparameters
I.1 Off-Policy
@@ I.3 Unsupervised RL
Appendix J Learning Curve
@@
Appendix K Full Result